OpenStudy (anonymous):

if a = 2 cm and r =3.4 cm , find the area of the regular pentagon

OpenStudy (anonymous):

\[Area = \frac{1}{2} \times r \times Perimeter\]

OpenStudy (anonymous):

Here: \[Perimeter = 5a = 5 \times 2 = 10\] \(r = 3.4\) Now plug in the values and find Area..